solve the system y-3z=-12 -2+y+2z=5 2x+3z=7
Answer by TimothyLamb(4379)

y-3z=-12 -2+y+2z=5 2x+3z=7
---
put the system of linear equations into standard form
---
y - 3z = -12
-2 + y + 2z = 5
2x + 3z = 7
---
y - 3z = -12
y + 2z = 7
2x + 3z = 7
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = -2.2
y = -0.6
z = 3.8
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php